OpenStudy (mathmath333):

Find the chance of throwing at least one ace in a simple throw of two dice ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ Find the chance of throwing at least one ace in a }\hspace{.33em}\\~\\ & \normalsize \text{ simple throw of two dice ?}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (jango_in_dtown):

ace is related to card problem.... how is ace coming in dice??

OpenStudy (zarkon):

ace means one

OpenStudy (mathmath333):

i m also confused with meaning of ace ,is it 7/36

OpenStudy (zarkon):

"The word "ace" comes from the Old French word as (from Latin 'as') meaning 'a unit', from the name of a small Roman coin. It originally meant the side of a die with only one mark, before it was a term for a playing card"

OpenStudy (jango_in_dtown):

I can provide the solution

OpenStudy (mathmath333):

(1,1)(1,2)(1,3)(1,4)(1,5)(1,6) is it 6/36=1/6

OpenStudy (zarkon):

no

OpenStudy (mathmath333):

u said ace =1 the numerator =6 [(1,1)(1,2)(1,3)(1,4)(1,5)(1,6)]

OpenStudy (zarkon):

(1,1)(1,2)(1,3)(1,4)(1,5)(1,6)(2,1)(3,1)(4,1)(5,1)(6,1)

OpenStudy (jango_in_dtown):

See let A denote the statement "no ace in 1st die". and B denote the event "no ace in second die" Then required probability is 1-P(AB) . Now in the first die, the probability of no ace=5/6 .also in the second die the probability of no ace=5/6. hence required probability =1-P(A)P(B) since the events are independent=1-(5/6)^2=11/36

OpenStudy (mathmath333):

oh ur right

OpenStudy (zarkon):

you could also use P(1 on die 1 or 1 on die 2) \[=\frac{1}{6}+\frac{1}{6}-\frac{1}{36}=\frac{11}{36}\]